OpenStudy (anonymous):

A rock is dropped from a seacliff and the sound of it striking the ocean is heard 3.4 seconds later. If the speed of sound is 340 m/s, how high is the cliff? I think the answer is 46 m but i don't get how it is that answer?

OpenStudy (anonymous):

\[h=ut + at^2/2\] Since the Stone is Dropped u = 0 \[h = at^2/2\] here a=g

OpenStudy (anonymous):

\[t _{1} + t _{2} =3.4\] \[h= 340 t _{2} \] \[h = \frac{ g(3.4-t _{2})^2 }{ 2 }\] \[ h = \frac{ g(3.4-\frac{ h }{ 340 })^2 }{ 2 }\]